OpenStudy (anonymous):

how do we find the fifth root of .0076 using log. Pls explain in full steps!!!!!!!!!

OpenStudy (anonymous):

Ok let y=0.0076^(1/5) take log on both sides log y=log(0.0076^(1/5)) now log (a^b)=b log a so log y=log(0.0076)/5 find log 0.0076 from log table then divide by 5. then take anti log on both sides to get y

OpenStudy (anonymous):

dude the book i was goin through gave a value of 1/5(3.8808){3 has a bar over it!!!!} which suddenly became equal to 1/5(5 + 2.8808){5 has a bar here}!!!!!!! How did dis happen

OpenStudy (anonymous):

since log 0.0076 is negative we cant directly take anti log so i'll take a random negative number and find anti log suppose we have -1.2 then add 2 and subtract 2 so -1.2 + 2 -2 -2 +0.8 find anti log (0.8) and then multiply by 10 ^-2

OpenStudy (anonymous):

ditch the bar and all

OpenStudy (anonymous):

bar = a - sign but one cannot substitute the bar for '-' sign so bar(-) 3.8808 = bar (-)5 +2.8808

OpenStudy (anonymous):

(-) 3 +0.8808 = (-) 5 +2 +0.8808